# UAV-VLA: Vision-Language-Action System for Large Scale Aerial Mission Generation

### v1

@@ Abstract
The UAV-VLA (Visual-Language-Action) system is a tool designed to facilitate communication with aerial robots. By integrating satellite imagery processing with the Visual Language Model (VLM) and the powerful capabilities of GPT, UAV-VLA enables users to generate general flight paths-and-action plans through simple text requests. This system leverages the rich contextual information provided by satellite images, allowing for enhanced decision-making and mission planning. The combination of visual analysis by VLM and natural language processing by GPT can provide the user with the path-and-action set, making aerial operations more efficient and accessible. The newly developed method showed the difference in the length of the created trajectory in 22% and the mean error in finding the objects of interest on a map in 34.22 m by Euclidean distance in the K-Nearest Neighbors (KNN) approach. The code is available here: https://github.com/sautenich/uav-vla
@@
VLA, VLM, LLM-agents, VLM-agents, UAV, Navigation, Drone, Path Planning.
@@ I Introduction
Many approaches to VLA and VLN necessitate extensive datasets that contain language instructions paired with sequences representing the agent behavior in the environment. These models are typically restricted to the specific environments utilized during their training, lacking the ability to generalize to novel contexts. Furthermore, they exhibit limited understanding of the global scale and do not possess a comprehensive representation of the surrounding environment. Our research emphasizes the development of systems capable of generating path plans and executing actions solely based on linguistic instructions and open satellite data using only zero-shot capabilities of the powerful models without any model training.
@@
  • We present a large-scale Vision-Language Action (VLA) system that generates complete path-action sets from a single text-based mission request, integrating textual inputs with satellite images.
  • We introduce the nano benchmark named UAV-VLPA-nano-30 aimed at fast measurements of the tasks solutions made by Vision Language Action systems at global scale.
  • We validate our system through the experiments on UAV-VLPA-nano-30, demonstrating performance comparable to human-level path and action generation.
@@ III-A Satellite Images And Metadata description
TABLE I: The specification of the UAV-VLPA-nano-30 benchmark
Name Samples
Total
Length,
km
Average
Length,
km
Place
UAV-VLPA-nano-30 30 63.89 2.13 United States of America
@@ III-B Manual Flight Plan Generation
The operator manually created flight plans for all 30 images in 35 minutes. An example of the Mission Planner environment with the violet square and home position is shown in Fig. 3. The total length of the UAV-VLPA-nano-30 benchmark samples and the average length created by the operator using the mission planner are listed in Table I.
@@ V-A Evaluation Metrics
The evaluation metric considers two aspects: the total length of the generated path and the error between each system-assigned point and the corresponding point in the human-generated trajectory (ground truth).
To compute the error, three methods were used to compare the system-generated and ground-truth trajectories. The Sequential Method aligns points step-by-step in their respective order, providing a measure of sequential similarity but prone to cumulative errors over longer trajectories.
Dynamic Time Warping (DTW) [21] enables non-linear alignment by adjusting the trajectories through stretching or compressing sections, effectively measuring path similarity without strict sequence matching.
K-Nearest Neighbors (KNN) matches each system-generated point to the nearest point in the ground truth based on spatial proximity, offering a general measure of accuracy without considering the order of points.
@@ VI Experimental Results
The newly developed system has shown a general trajectory length of 77.74 km on the benchmark UAV-VLA-nano-30, which is 13.85 km, or 21.6%, longer than the ground-truth trajectory created by the experienced UAV pilot (Section III). In 7 out of 30 cases, or 23% of the cases, the UAV-VLA generated a trajectory path that was even shorter as can be seen in Fig. 5.
@@
As shown in Table II, the sequential RMSE exhibited the largest mean error of 409.54 m per trajectory, which was expected due to its strict reliance on the sequential order of points. The Dynamic Time Warping (DTW) method demonstrated a reduced mean error of 307.27 m, highlighting its ability to account for temporal variations more effectively. The K-Nearest Neighbors (KNN) method resulted in the smallest mean error, as it disregards the sequence entirely and focuses solely on the spatial proximity of points.
TABLE II: Comparison of RMSE Metrics for Different Methods
@@
VII Discussion
This paper presents a novel approach for UAV Mission Generation on a global scale, enhancing flexibility and accuracy in mission planning. By addressing the limitations of traditional manual methods, this approach proves valuable in scenarios where manual intervention is inefficient. The main contributions of this work include:
  • The benchmark UAV-VLPA-nano-30, providing a standardized framework for evaluating global-scale path planning techniques.
  • A method to interpret natural language requests into actionable flight paths, generating paths only 21.6% longer than human-created ones, showcasing its efficiency.
  • A new task for UAVs: language-based path planning, enabling autonomous execution of mission plans from natural language inputs.
This approach simplifies human-UAV interaction by enabling direct communication via natural language, eliminating intermediate devices. Additionally, it lays the groundwork for robot-robot interaction, allowing autonomous mission generation between robots. This innovation paves the way for seamless collaboration between UAVs, humans, and other robots in diverse environments.
@@ VIII Future Work
Future work will focus on creating a specialized dataset for training models in satellite map-based path planning. This dataset will enhance model precision and efficiency in mission generation across various UAV applications. Additionally, we aim to develop an end-to-end model that autonomously generates mission plans from high-level goals, integrating action generation, path planning, and decision-making into a unified framework. This will represent a significant step towards achieving fully autonomous UAV mission planning adaptable to diverse environments and objectives.

### v2

@@ Abstract
The UAV-VLA (Visual-Language-Action) system is a tool designed to facilitate communication with aerial robots. By integrating satellite imagery processing with the Visual Language Model (VLM) and the powerful capabilities of GPT, UAV-VLA enables users to generate general flight paths-and-action plans through simple text requests. This system leverages the rich contextual information provided by satellite images, allowing for enhanced decision-making and mission planning. The combination of visual analysis by VLM and natural language processing by GPT can provide the user with the path-and-action set, making aerial operations more efficient and accessible. The newly developed method showed the difference in the length of the created trajectory in 22% and the mean error in finding the objects of interest on a map in 34.22 m by Euclidean distance in the K-Nearest Neighbors (KNN) approach. Additionally, the UAV-VLA system generates all flight plans in just 5 minutes and 24 seconds, making it 6.5 times faster than an experienced human operator. The code is available here: https://github.com/sautenich/uav-vla
@@
VLA; VLM; LLM-agents; VLM-agents; UAV; Navigation; Drone; Path Planning.
@@ I Introduction
Many VLA and VLN approaches require large datasets with paired language instructions and agent behavior, but struggle to generalize to new environments and lack a global-scale understanding. Our research focuses on developing systems that generate path plans and execute actions based solely on linguistic instructions and open satellite data, leveraging zero-shot capabilities of powerful models without additional training.
@@
  • We present a large-scale Vision-Language-Action (VLA) system that generates path-action sets from text-based mission requests using satellite images.
  • We introduce the UAV-VLPA-nano-30 benchmark for evaluating VLA systems at a global scale.
  • We validate our system on UAV-VLPA-nano-30, showing performance comparable to human-level path and action generation.
@@ III-B Manual Flight Plan Generation
The operator manually created flight plans for all 30 images in 35 minutes. An example of the Mission Planner environment with the violet square and home position is shown in Fig. 3. The total length of the UAV-VLPA-nano-30 benchmark samples was 63.89 km and the average length was 2.13 km created by the operator using the Mission Planner.
@@ V-A Evaluation Metrics
The evaluation metric considers two aspects: the total length of the path and the error between system-generated and human-generated (ground truth) trajectories.
Three methods were used to compute the error: The Sequential Method aligns points step-by-step but is prone to cumulative errors. Dynamic Time Warping (DTW) [21] allows non-linear alignment to measure path similarity. K-Nearest Neighbors (KNN) matches system points to the nearest ground-truth points, offering a general accuracy measure without considering order.
@@ VI Experimental Results
The newly developed system has shown a general trajectory length of 77.74 km on the benchmark UAV-VLA-nano-30, which is 13.85 km, or 21.6%, longer than the ground-truth trajectory created by the experienced UAV pilot (Sec. III-B). In 7 out of 30 cases, or 23% of the cases, the UAV-VLA generated a trajectory path that was even shorter as can be seen in Fig. 5.
@@
As shown in Table I, the sequential RMSE exhibited the largest mean error of 409.54 m per trajectory, which was expected due to its strict reliance on the sequential order of points. The Dynamic Time Warping (DTW) method demonstrated a reduced mean error of 307.27 m, highlighting its ability to account for temporal variations more effectively. The K-Nearest Neighbors (KNN) method resulted in the smallest mean error, as it disregards the sequence entirely and focuses solely on the spatial proximity of points.
TABLE I: Comparison of RMSE Metrics for Different Methods
@@
VII Conclusion
This paper presents a global-scale UAV mission generation approach, enhancing flexibility and accuracy in mission planning. By overcoming the limitations of traditional manual methods, this approach is highly beneficial in scenarios where human intervention is inefficient. Key contributions include:
  • The UAV-VLPA-nano-30 benchmark, providing a standardized framework for evaluating and comparing global path planning techniques.
  • A method for interpreting natural language requests into flight paths, generating paths that are 21.6% longer but 6.5 times faster than human-created ones, demonstrating significant efficiency.
  • A new task for UAVs: language-based path planning, enabling UAVs to autonomously execute mission plans from natural language inputs.
This method simplifies human-UAV interaction by enabling direct natural language communication and lays the foundation for seamless robot-robot collaboration in dynamic environments.
@@ VIII Future Work
Future work will focus on creating a specialized dataset for satellite map-based path planning to improve model precision and efficiency in UAV mission generation. Additionally, we aim to develop an end-to-end model that autonomously generates mission plans from high-level goals, integrating action generation, path planning, and decision-making for fully autonomous UAV mission planning across diverse environments.